Reference sheet for natbib usage

(Describing version 6.8 from 1998/02/19)

For a more detailed description of the natbib package, L^{A}TEX the source file natbib.dtx.

Overview

The natbib package is a reimplementation of the L^{A}TEX \cite command, to work with both author–year and numerical citations. It is compatible with the standard bibliographic style files, such as plain.bst, as

well as with those for harvard, apalike, chicago, astron, authordate, and of course natbib.

Loading

Load with \usepackage[options]{natbib}. See list of options at the end.

Basic commands

The natbib package has two basic citation commands, \citet and \citep for textual and parenthetical citations, respectively. There also exist the starred versions \citet* and \citep* that print the full author list, and not just the abbreviated one. All of these may take one or two optional arguments to add some text before and after the citation.

\citet{jon90}  ⇒ Jones et al. (1990)
\citet[chap.~2]{jon90}  ⇒ Jones et al. (1990, chap. 2)
\citep{jon90}  ⇒ (Jones et al., 1990)
\citep[chap.~2]{jon90}  ⇒ (Jones et al., 1990, chap. 2)
\citep[see][]{jon90}  ⇒ (see Jones et al., 1990)
\citep[see][chap.~2]{jon90} ⇒ (see Jones et al., 1990, chap. 2)
\citet*{jon90}  ⇒ Jones, Baker, and Williams (1990)
\citep*{jon90}  ⇒ (Jones, Baker, and Williams, 1990)

Multiple citations

Multiple citations may be made as usual, by including more than one citation key in the \cite command argument.

\citet{jon90,jam91}  ⇒ Jones et al. (1990); James et al. (1991)
\citep{jon90,jam91}  ⇒ (Jones et al., 1990; James et al. 1991)
\citep{jon90,jon91}  ⇒ (Jones et al., 1990, 1991)

      \citep{jon90a,jon90b} ⇒ (Jones et al., 1990a,b)

Numerical mode

These examples are for author–year citation mode. In numerical mode, the results are different.

\citet{jon90}  ⇒ Jones et al. [21]

\citet[chap.~2]{jon90}  ⇒ Jones et al. [21, chap. 2]
\citep{jon90}  ⇒ [21]
\citep[chap.~2]{jon90}  ⇒ [21, chap. 2]
\citep[see][]{jon90}  ⇒ [see 21]
\citep[see][chap.~2]{jon90} ⇒ [see 21, chap. 2]
\citep{jon90a,jon90b}  ⇒ [21, 32]

Suppressed parentheses

As an alternative form of citation, \citealt is the same as \citet but without any parentheses. Similarly, \citealp is \citep with the parentheses turned off. Multiple references, notes, and the starred variants also exist.

\citealt{jon90}  ⇒ Jones et al. 1990

\citealt*{jon90}  ⇒ Jones, Baker, and Williams 1990
\citealp{jon90}  ⇒ Jones et al., 1990
\citealp*{jon90}  ⇒ Jones, Baker, and Williams, 1990
\citealp{jon90,jam91}  ⇒ Jones et al., 1990; James et al., 1991

\citealp[pg.~32]{jon90} ⇒ Jones et al., 1990, pg. 32
\citetext{priv.\ comm.} ⇒ (priv. comm.)

The \citetext command allows arbitrary text to be placed in the current citation parentheses. This may be used in combination with \citealp.

Partial citations

In author–year schemes, it is sometimes desirable to be able to refer to the authors without the year, or vice versa. This is provided with the extra commands

\citeauthor{jon90} ⇒ Jones et al.

\citeauthor*{jon90} ⇒ Jones, Baker, and Williams

\citeyear{jon90}  ⇒ 1990
\citeyearpar{jon90} ⇒ (1990)

Selecting citation style and punctuation

Use the command \bibpunct with one optional and 6 mandatory arguments:

1. the opening bracket symbol, default = (

2. the closing bracket symbol, default = )

3. the punctuation between multiple citations, default = ;

4. the letter ‘n’ for numerical style, or ‘s’ for numerical superscript style, any other letter for author–year, default = author–year;

5. the punctuation that comes between the author names and the year

6. the punctuation that comes between years or numbers when common author lists are suppressed

(default = ,); The optional argument is the character preceding a post-note, default is a comma.

Example 1, \bibpunct{[}{]}{,}{a}{}{;} changes the output of

     \citep{jon90,jon91,jam92}

into [Jones et al. 1990; 1991, James et al. 1992].

Example 2, \bibpunct[;]{(}{)}{,}{a}{}{;} changes the output of
     \citep[and references therein]{jon90}

into (Jones et al. 1990; and references therein).

Other formatting options

Redefine \bibsection to the desired sectioning command for introducing the list of references. This is

normally \section* or \chapter*.

Redefine \bibfont to be a font declaration, e.g. \small to apply to the list of references. The indentation after the first line of each reference is given by \bibhang; change this with the \setlength command. The vertical spacing between references is set by \bibsep; change this with the \setlength command.

Automatic indexing of citations

If one wishes to have the citations entered in the .idx indexing file, it is only necessary to issue \citeindextrue at any point in the document. All following \cite commands, of all variations, then insert the corresponding entry to that file. With \citeindexfalse, these entries will no longer be made.

Sorting and compressing citations

Do not use the cite package with natbib; rather use one of the options sort or sort&compress. These also work with author–year citations, making multiple citations appear in their order in the reference list.

Long author list on first citation

Use option longlistfirst to have first citation automatically give the full list of authors. Suppress this for certain citations with \shortcites{key-list}, given before the first citation.

Replacement bibliography styles

I provide three new .bst files to replace the standard L^{A}TEX numerical ones:

     plainnat.bst     abbrvnat.bst     unsrtnat.bst

Local configuration

Any local recoding or definitions can be put in natbib.cfg which is read in after the main package file.

Options that can be added to \usepackage

round (default) for round parentheses; square for square brackets; curly for curly braces; angle for angle brackets; colon (default) to separate multiple citations with colons; comma to use commas as separaters; authoryear (default) for author–year citations; numbers for numerical citations; super for superscripted numerical citations, as in Nature; sort orders multiple citations into the sequence in which they appear in the list of references; sort&compress as sort but in addition multiple numerical citations are compressed if possible (as 3–6, 15); longnamesfirst makes the first citation of any reference the equivalent of the starred variant (full author

list) and subsequent citations normal (abbreviated list);

sectionbib redefines \thebibliography to issue \section* instead of \chapter*; valid only for classes

with a \chapter command; to be used with the chapterbib package.